# PAC-Bayes Analysis for Recalibration in Classification

### v1

@@ Abstract
Nonparametric estimation with binning is widely employed in the calibration error evaluation and the recalibration of machine learning models. Recently, theoretical analyses of the bias induced by this estimation approach have been actively pursued; however, the understanding of the generalization of the calibration error to unknown data remains limited. In addition, although many recalibration algorithms have been proposed, their generalization performance lacks theoretical guarantees. To address this problem, we conduct a generalization analysis of the calibration error under the probably approximately correct (PAC) Bayes framework. This approach enables us to derive a first optimizable upper bound for the generalization error in the calibration context. We then propose a generalization-aware recalibration algorithm based on our generalization theory. Numerical experiments show that our algorithm improves the Gaussian-process-based recalibration performance on various benchmark datasets and models.
@@ 1 Introduction
Increasing the reliability of machine learning models is crucial in risk-sensitive applications such as autonomous driving [4]. Recently, the concept of calibration has become a significant measure of reliability, especially in classification tasks. In this context, the calibration performance is evaluated by how well predictive probabilities provided by our model align with the actual frequency of true labels. A close correspondence between them indicates that the model is well calibrated [6, 33]. To evaluate the calibration performance, the calibration error (CE) [11, 27, 9] or the top-label CE (TCE) [18, 9] is often used. This evaluates the disparity between the predicted probability of a model and the conditional probability of the label frequency given by the model prediction. Analytically computing the CE or TCE is, however, challenging because the conditional probabilities of the labels are intractable. A common approach to addressing this issue is constructing its estimator called the expected calibration error (ECE) by binning [35].
If the ECE is small, we consider the model well calibrated and its predictions highly reliable. Conversely, if the model is poorly calibrated, we correct the predictive probability by a post processing method called recalibration [10, 35] using a recalibration dataset that is independent of the train and test datasets. In the recalibration context, it is significant to theoretically guarantee that the recalibrated model can indeed achieve good calibration in terms of the ECE for unknown data. However, no studies have been conducted from this perspective to the best of our knowledge, although there has been extensive research into examining the estimation bias of the ECE from the CE [18, 11, 9, 29] and proposing specific algorithms for recalibration [24, 36, 23, 17, 10, 32, 38]. Additionally, no research has been conducted to analyze the generalization nature of the ECE in the case of nonrecalibration. This implies that there is no theoretical guarantee that a model achieving a small ECE on training data also achieves a small ECE on unknown data. In other words, regardless of recalibration, the generalization theory under the ECE is not yet established.
@@
In this study, we construct a novel generalization theory for the ECE before and after recalibration using probably approximately correct (PAC) Bayes theory [21, 1], which is the common method of deriving generalization bounds across a wide range of models. Since PAC-Bayes theory can analyze generalization performance in an algorithm-dependent manner, the obtained upper bound is often optimizable and useful for deriving new generalization-aware algorithms. In applying PAC-Bayes analysis to the ECE, we face the following two challenges: (i) Owing to the nonparametric estimation of conditional probabilities, the ECE computed on the test dataset is not a sum of independent and identically distributed (i.i.d.) random variables and thus cannot be simply applied using the existing PAC-Bayes bounds derived under the i.i.d. assumption, and (ii) some bins may not contain samples because binning divides the probability interval $[0,1]$ into equal widths, making it difficult to apply the concentration inequality used in the PAC-Bayes bound derivation.
Our main contribution is the establishment of a generalization error bound in the ECE, including recalibration scenarios, by deriving new concentration inequalities that address the above challenges. This theory also enables the analysis of the bias between the ECE and the CE or the TCE through PAC-Bayes analysis. The results reveal the convergence rate of the generalization error and bias, $\mathcal{O}(1/n^{1/3})$, regardless of whether recalibration is performed, where $n$ is the number of data points for training or recalibration. Furthermore, our theory makes it possible to analyze the ECE in both binary and multiclass classifications. We then show that although the ECE can only capture the calibration performance for the top-$1$ predictive probability, it is unaffected by the curse of dimensionality.
Furthermore, we formulate a novel recalibration algorithm by leveraging our PAC-Bayes bound, which is expected to improve the generalization performance in terms of ECE and reduce bias in CE or TCE. Our method can be regarded as an extension of the recently proposed recalibration method based on Gaussian processes [32], incorporating flexible a Kullback–Leibler (KL) regularization with control parameters and calibration errors as a recalibration loss. Numerical experiments confirm a correlation between the KL terms in our bounds and the generalization performance, show that our method can improve GP-based recalibration, and reveal the instability of ECE-based calibration performance evaluation due to the slow convergence rate of $\mathcal{O}(1/n^{1/3})$.
@@ 2 Preliminaries
In this section, we summarize the basic notations, problem setup (Section 2.1), calibration metric (Section 2.2), and the generalization notion in recalibration (Section 2.3).
@@ 2.1 Notations and problem setting
For a random variable denoted in capital letters, we express its realization with corresponding lowercase letters. Let $P(X)$ denote a probability of $X$, and let $P(Y|X)$ represent the conditional probability of $Y$ given $X$. We express the expectation of a random variable $X$ as $\mathbb{E}_{X}$. Let $\mathrm{KL}(P\|Q)$ be the Kullback–Leibler (KL) divergence of $P$ from $Q$, where $Q$ is a probability distribution and $P$ is absolutely continuous with respect to $Q$.
Let $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ be the domain of data, where $\mathcal{X}$ and $\mathcal{Y}$ are input and label spaces, respectively. Suppose $\mathcal{D}$ represents an unknown data distribution, and let $S_{\mathrm{tr}}\coloneqq\{Z_{m}\}_{m=1}^{n_{\mathrm{tr}}}$ denote the training dataset consisting of $n_{\mathrm{tr}}$ samples drawn independently and identically distributed (i.i.d.) drawn from $\mathcal{D}$. We also define the test dataset comprising $n_{\mathrm{te}}$ samples as $S_{\mathrm{te}}\sim\mathcal{D}^{n_{\mathrm{te}}}$. Let $f_{w}:\mathcal{X}\to\mathcal{Y}$ be our parametric predictor and we denote its parameters as $w\in\mathcal{W}\subset\mathbb{R}^{d}$. We evaluate the performance of the trained predictor $f_{w}$ using the loss function $l:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$, where $l(y,f_{w}(x))$ denotes the loss incurred by the prediction $f_{w}(x)$ for the target $y$. We define the training loss as $\hat{L}(w,S_{\mathrm{tr}})\coloneqq\frac{1}{n_{\mathrm{tr}}}\sum_{(X,Y)\in S_{\mathrm{tr}}}l(Y,f_{w}(X))$ and the (expected) test loss as $L(w,\mathcal{D})\coloneqq\mathbb{E}_{S_{\mathrm{te}}}\frac{1}{n_{\mathrm{te}}}\sum_{(X,Y)\in S_{\mathrm{te}}}l(Y,f_{w}(X))$. Under this setting, the generalization error for $l$ is defined as $\mathrm{gen}(f_{w},S_{\mathrm{tr}},l)\coloneqq L(w,\mathcal{D})-\hat{L}(w,S_{\mathrm{tr}})$. The following theorem is the most basic form of the PAC-Bayes generalization bounds [1].
@@
2.2 Calibration metrics
Calibration for binary classification:
Let $\mathcal{Y}=\{0,1\}$ and let $f_{w}:\mathcal{X}\to[0,1]$ be a model, where the output corresponds to the model’s confidence that the label is $1$. In the context of calibration, it is expected that not only $f_{w}$ has high accuracy, but also its predictive probability aligns well in the actual label frequency $P(Y|f_{w}(X))$. A model is well calibrated [33, 12] when $P(Y|f_{w}(X))=f_{w}(X)$ holds; however, it is difficult to confirm this. Therefore, an alternative metric called as the calibration error (CE) [33, 12] is typically used. Given $W=w$, the CE is defined as
Unfortunately, evaluating the CE is still infeasible because $\mathbb{E}[Y|f_{w}(X)]$ is intractable. To resolve this issue, we evaluate the calibration performance by constructing the estimator of the CE. The widely used estimator is the expected calibration error (ECE), where binning is used to estimate $\mathbb{E}[Y|f_{w}(X)]$ [10, 35, 36]. The ECE is calculated by partitioning the range of the predictive probability $[0,1]$ into $B$ intervals =$\mathcal{I}=\{I_{i}\}_{i=1}^{B}$ (called bins) and averaging within each bin using an evaluation dataset $S_{e}\coloneqq\{z_{m}\}_{m=1}^{n_{e}}\in\mathcal{Z}^{n_{e}}$, where we assume $n_{e}\geq 2B$. That is, the ECE is defined as
Calibration for multiclass classification:
Let the label space be $Y\in\mathcal{Y}\coloneqq\{0,\dots,K-1\}$, where $K\in\mathbb{N}$ is the number of classes. We then define the one-hot encoding of the label and its domain as $\hat{Y}\in\hat{\mathcal{Y}}\subset\mathbb{R}^{K}$. Let $f_{w}:\mathcal{X}\to\Delta^{K}$ be a probabilistic classifier, where $\Delta^{K}$ represents the $K-1$ dimensional simplex. For example, such a simplex is obtained by the final softmax layer in neural networks. Under these settings, $f_{w}$ predicts the label by $C\coloneqq\mathrm{argmax}_{k}f_{w}(X)_{k}$, where $f_{w}(X)_{k}$ is the $k$-th dimension of $f_{w}(x)$, which represents the model’s confidence of the label $k\in K$.
A frequently used measure in the above setup is the top-label calibration error (TCE) [18, 9], which is defined as follow with respect to the highest prediction probability output by $f_{w}$:
Similarly to $\mathrm{CE}(f_{w})$, the estimator of the TCE can be obtained by binning to $f_{w}(X)_{C}$. That is, we divide the range of the predictive probability of $f_{w}(X)_{C}$ into $B$ bins of the same width, and then calculate the following ECE: $\mathrm{ECE}(f_{w},S_{e})\coloneqq\sum_{i=1}^{B}p_{i}|\bar{f}_{i,S_{e}}-\bar{p}_{i,S_{e}}|$, where $|I_{i}|\coloneqq\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}$, $p_{i}\coloneqq|I_{i}|/n_{e}$, $\bar{f}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}f_{w}(x_{m})$, and $\bar{p}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}\mathbbm{1}_{y_{m}=C}$. For brevity, we refer to the TCE estimator in both binary and multiclass classifications constructed by binning as the ECE.
2.3 Generalization for recalibration analysis
The trained predictor $f_{w}(x)$ can be poorly calibrated [10, 18]. One common approach to address this problem involves recalibrating $f_{w}(x)$ by post processing using the parametric function  [^1] In the binary case, given $x$, we use $(f_{w}(x),1-f_{w}(x))\in\Delta^{2}$ as the input to $\eta_{V}(\cdot)$; see Appendix B for details. $\eta_{V}:\Delta^{K}\to\Delta^{K}$, where $V\in\mathcal{V}\subset\mathbb{R}^{d^{\prime}}$ is the parameter learned by the recalibration dataset $S_{\mathrm{re}}\sim\mathcal{D}^{n_{\mathrm{re}}}$ at fixed $w$ and $S_{\mathrm{re}}$ is independent of $S_{\mathrm{tr}}$. In this procedure, the overall dataset is split into training data $S_{\mathrm{tr}}$ for learning $W$, the recalibration data $S_{\mathrm{re}}$ for learning $V$, and the test data $S_{\mathrm{te}}$ for the evaluation of the ECE. We define the recalibrated model as $\eta_{v}(f_{w})$.
We expect $\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{te}})$ to be sufficiently small after recalibration. Therefore, it is important to theoretically explore the conditions for achieving an equally small $\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{te}})$ for the unknown test data $S_{\mathrm{te}}$ when $\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})$ is sufficiently small. To this end, we study the following error:
$\mathrm{gen}(\eta_{v}(f_{w}),S_{\mathrm{re}},\mathrm{ECE})\coloneqq|\mathbb{E}_{S_{\mathrm{te}}}\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})|.$ (5)
We refer to this quantity as the generalization error of the ECE under $S_{\mathrm{re}}$.
Given that the ECE is an estimator of the CE or the TCE obtained by binning, it is important to evaluate how recalibration affects the bias between them expressed as
$\mathrm{Bias}(\eta_{v}(f_{w}),S_{\mathrm{re}},\mathrm{TCE})\coloneqq|\mathrm{TCE}(\eta_{v}(f_{w}))-\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})|.$ (6)
We refer to this as the estimation bias of the ECE under $S_{\mathrm{re}}$. Similarly, we define the estimation bias of CE as $\mathrm{Bias}(\eta_{v}(f_{w}),S_{\mathrm{re}},\mathrm{CE})\coloneqq|\mathrm{CE}(\eta_{v}(f_{w}))-\mathrm{ECE}(\eta_{v}(f_{w}),S_{\mathrm{re}})|$ for binary classification.
3 PAC-Bayes bounds for ECE without recalibration
Our goal is to provide a comprehensive analysis of Eqs. (5) and (6) by utilizing the PAC-Bayes theory. Before showing our main results for them, we present our PAC-Bayes analysis for the following two quantities: $\mathrm{gen}(f_{w},S_{\mathrm{tr}},\mathrm{ECE})\coloneqq|\mathbb{E}_{S_{\mathrm{te}}}\mathrm{ECE}(f_{w},S_{\mathrm{te}})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|$ and $\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{CE})\coloneqq|\mathrm{CE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|$, for which we do not conduct recalibration. The results of this analysis are crucial for our recalibration analysis in Section 4 and provide a rigorous understanding of the generalization and estimation bias in the ECE in the absence of recalibration.
3.1 Generalization error analysis for calibration

### v2

@@ Abstract
Nonparametric estimation using uniform-width binning is a standard approach for evaluating the calibration performance of machine learning models. However, existing theoretical analyses of the bias induced by binning are limited to binary classification, creating a significant gap with practical applications such as multiclass classification. Additionally, many parametric recalibration algorithms lack theoretical guarantees for their generalization performance. To address these issues, we conduct a generalization analysis of calibration error using the probably approximately correct Bayes framework. This approach enables us to derive the first optimizable upper bound for generalization error in the calibration context. On the basis of our theory, we propose a generalization-aware recalibration algorithm. Numerical experiments show that our algorithm enhances the performance of Gaussian process-based recalibration across various benchmark datasets and models.
Machine Learning, ICML, Calibration, PAC-Bayes
@@ 1 Introduction
Increasing the reliability of machine learning models is crucial in risk-sensitive applications such as autonomous driving (Chen et al., 2015). Recently, the concept of calibration has become a significant measure of reliability, especially in classification tasks. In this context, the calibration performance is evaluated by how well predictive probabilities provided by our model align with the actual frequency of true labels. A close correspondence between them indicates that the model is well-calibrated (Dawid, 1982; Widmann et al., 2021). To evaluate the calibration performance, a calibration error (Gupta and Ramdas, 2021; Roelofs et al., 2022) such as the top-label calibration error (TCE) (Kumar et al., 2019; Gruber and Buettner, 2022) is often used. This evaluates the disparity between the predicted probability of a model and the conditional probability of the label frequency given by the model prediction. Analytically computing the TCE is, however, challenging because the conditional probabilities of the labels are intractable. Among various methods proposed to address this issue, constructing its estimator called the expected calibration error (ECE) using uniform-width binning (UWB) (Zadrozny and Elkan, 2001; Naeini et al., 2015) is one of the most widely adopted. This study centers on this approach.
If the ECE is small, we consider the model well-calibrated and its predictions highly reliable. Unfortunately, as has already been made evident in recent studies, models such as neural networks are not necessarily well-calibrated (Guo et al., 2017). If a model shows poor calibration performance, it is common to apply a post-processing technique known as recalibration (Guo et al., 2017; Zadrozny and Elkan, 2001). This technique adjusts predicted probabilities using a recalibration function—a parametric function trained separately from the original model—on a dataset independent of both training and test data. Recalibration aims to train a separate function that, when composed with the original predictor, yields a low ECE. Numerous methodologies have been proposed recently, including temperature scaling (Guo et al., 2017) and recalibration based on variational inference (VI) with a Gaussian process (GP) (Wenger et al., 2020).
Given that the ECE is an estimator of the TCE, it is important for reliable uncertainty evaluation to understand the extent of bias introduced between them before and after recalibration. Nevertheless, the theoretical understanding of this remains limited. In the context of bias analysis before recalibration, many studies have focused exclusively on binary classification and have shown that the number of bins used in binning significantly affects both the estimated ECE and the bias (Gupta and Ramdas, 2021; Sun et al., 2023; Futami and Fujisawa, 2024). Such insights remain unclear in the multiclass classification context. Moreover, regularization techniques have been proposed to reduce the ECE on training data (Kumar et al., 2018; Popordanoska et al., 2022; Wang et al., 2021a). However, a low ECE on training data does not guarantee a similar low TCE. Verifying this requires a theoretical analysis of the gap between these, which, to our knowledge, has not yet been explored.
This situation remains the same even after recalibration. Moreover, in many recalibration methods, postprocessing is performed to achieve a low ECE only on the recalibration data. However, in practice, it is desirable to perform recalibration such that the ECE evaluated on test data is reduced—that is, to enhance the model’s generalization performance from the ECE perspective through recalibration. Although existing studies have confirmed this point only through numerical experiments (Platt, 1999; Zadrozny and Elkan, 2002; Kull et al., 2017; Naeini et al., 2015; Guo et al., 2017; Wenger et al., 2020), recalibration methods with theoretical guarantees for improving generalization performance in terms of ECE have not been explored thus far.
@@
Given this background, we conduct a comprehensive analysis of the bias of the ECE in multiclass classification. To achieve this, we develop a statistical learning theory focusing on the ECE before and after recalibration by applying the probably approximately correct (PAC) Bayes theory (McAllester, 2003; Alquier et al., 2016), a general method for deriving generalization bounds across various models. Since PAC-Bayes theory allows for the algorithm-dependent analysis of generalization performance, the obtained upper bounds are often optimizable and useful for deriving new generalization-aware algorithms. In applying PAC-Bayes analysis to the ECE, we face the following two challenges: (i) Owing to the nonparametric estimation of conditional probabilities, the ECE computed on the test dataset is not a sum of independently and identically distributed (i.i.d.) random variables and thus cannot be applied using the existing PAC-Bayes bounds derived under the i.i.d. assumption, and (ii) some bins may not contain samples because UWB divides the probability interval $[0,1]$ into equal widths, making it difficult to apply the concentration inequality used in the PAC-Bayes bound derivation.
Our main contribution is a novel analysis of ECE in multiclass classification, which addresses the abovementioned limitations. We begin by presenting a decomposition of the bias into binning bias and finite-sample estimation bias. We then derive a novel concentration inequality that enables theoretical analysis in the binning-based ECE setting with UWB. This framework reveals that the bias in ECE estimation converges at a notably slow rate. Furthermore, by deriving the PAC-Bayes bound, we have successfully formulated a new generalization-aware recalibration algorithm, which is expected to improve the generalization performance of calibration and reduce the estimation bias in TCE. Numerical experiments confirm a correlation between the Kullback–Leibler (KL) regularization terms in our bounds and the generalization performance, show that our method can improve GP-based recalibration, and reveal the instability of ECE-based calibration performance evaluation due to the slow convergence rate. Finally, we discuss related work in light of our theoretical findings in Section 5, providing a view of how our contributions relate to existing literature.
@@ 2 Preliminaries
In this section, we summarize the basic notations, problem setup (Section 2.1), calibration metric (Section 2.2), and postprocessing by recalibration (Section 2.3).
@@ 2.1 Notations and Problem Setting
For a random variable denoted in capital letters, we express its realization with corresponding lowercase letters. Let $P(X)$ denote a probability distribution of $X$, and let $P(Y|X)$ represent the conditional probability distribution of $Y$ given $X$. We express the expectation of a random variable $X$ as $\mathbb{E}_{X}$. Let $\mathrm{KL}(P\|Q)$ be the KL divergence of $P$ from $Q$, where $Q$ is a probability distribution and $P$ is absolutely continuous with respect to $Q$.
Let $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$ be the domain of data, where $\mathcal{X}$ and $\mathcal{Y}$ are the input and label spaces, respectively. Let the label space be $\mathcal{Y}\coloneqq\{0,\dots,K-1\}$, where $K\in\mathbb{N}$ is the number of classes. For the label $Y\in\mathcal{Y}$, we define the one-hot encoding of the label $Y$ as $\mathbf{e}_{Y}\in\mathbb{R}^{K}$. Suppose $\mathcal{D}$ represents an unknown data distribution, and let $S_{\mathrm{tr}}\coloneqq\{(X_{m},Y_{m})\}_{m=1}^{n_{\mathrm{tr}}}$ denote the training dataset consisting of $n_{\mathrm{tr}}$ samples drawn i.i.d. from $\mathcal{D}$. We also define the test dataset comprising $n_{\mathrm{te}}$ samples drawn i.i.d. from $\mathcal{D}$. Let $f_{w}:\mathcal{X}\to\Delta^{K}$ be a probabilistic classifier, where $\Delta^{K}$ represents the $K-1$-dimensional simplex. This classifier is parameterized by $w\in\mathcal{W}\subset\mathbb{R}^{d}$. For example, such a simplex is obtained by the final softmax layer in neural networks. Under these settings, $f_{w}$ predicts the label from $C\coloneqq\mathrm{argmax}_{k}f_{w}(X)_{k}$, where $f_{w}(X)_{k}$ is the $k$-th dimension of $f_{w}(x)$, which represents the model’s confidence of the label $k\in K$.
We evaluate the performance characteristics of the trained predictor $f_{w}$, such as its accuracy using the classification loss function $l_{\mathrm{acc}}:\mathcal{Y}\times\Delta^{K}\to\mathbb{R}$, where $l_{\mathrm{acc}}(y,f_{w}(x))$ denotes the loss incurred by the prediction $f_{w}(x)$ for the target $y$. We define the training loss as $\hat{L}(w,S_{\mathrm{tr}})\coloneqq\frac{1}{n_{\mathrm{tr}}}\sum_{(X,Y)\in S_{\mathrm{tr}}}l_{\mathrm{acc}}(Y,f_{w}(X))$ and the expected loss as $L(w,\mathcal{D})\coloneqq\mathbb{E}_{(X,Y)}l_{\mathrm{acc}}(Y,f_{w}(X))$. Under this setting, one of the major purposes of the learning algorithms is to achieve the small generalization gap for loss function $l_{\mathrm{acc}}$, which is defined as $|L(w,\mathcal{D})-\hat{L}(w,S_{\mathrm{tr}})|$.
2.2 Calibration Metrics for Multiclass Classification
In the context of calibration, it is expected that not only $f_{w}$ has high accuracy but also its predictive probability aligns well in the actual label frequency $P(Y=k|f_{w}(x))$ for $k=0,\dots,K-1$. Hereinafter, given $(w,x)$, we treat $P(Y|f_{w}(x))$ as the $K$-dimensional vector by regarding its $k$-th element as $P(Y=k|f_{w}(x))$. Then, a model is well-calibrated (Widmann et al., 2021; Gupta et al., 2020) when $P(Y|f_{w}(X))=f_{w}(X)$ holds. However, it is difficult to confirm this in practice. Instead, a frequently used measure is the top-label calibration error (TCE) (Kumar et al., 2019; Gruber and Buettner, 2022), which uses the highest prediction probability in $f_{w}$:
Unfortunately, evaluating the TCE is still infeasible because $P(Y=C|f_{w}(X)_{C})$ is intractable. To resolve this issue, we evaluate the calibration performance by constructing the estimator of the TCE. The widely used estimator is the expected calibration error (ECE), where binning is used to estimate $P(Y=C|f_{w}(X)_{C})$ (Guo et al., 2017; Zadrozny and Elkan, 2001, 2002). The ECE is often computed using UWB, which partitions the predictive probability range $[0,1]$ into $B$ equal-width intervals =$\mathcal{I}=\{I_{i}\}_{i=1}^{B}$ (called bins) and averaging within each bin using an evaluation dataset $S_{e}\coloneqq\{z_{m}\}_{m=1}^{n_{e}}\in\mathcal{Z}^{n_{e}}$. For simplicity, we refer to UWB simply as binning. Accordingly, we define the ECE based on binning as follows:
$\mathrm{ECE}(f_{w},S_{e})\coloneqq\sum_{i=1}^{B}p_{i}|\bar{f}_{i,S_{e}}-\bar{p}_{i,S_{e}}|,$ (2)
where $|I_{i}|\coloneqq\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}$, $p_{i}\coloneqq\frac{|I_{i}|}{n_{e}}$, $\bar{f}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}f_{w}(x_{m})_{C}$, and $\bar{p}_{i,S_{e}}\coloneqq\frac{1}{|I_{i}|}\sum_{m=1}^{n_{e}}\mathbbm{1}_{f_{w}(x_{m})_{C}\in I_{i}}\mathbbm{1}_{y_{m}=C}$. Here, bins are set by dividing the interval $[0,1]$ into $B$ bins of the same width: $I_{1}=(0,1/B],I_{2}=(1/B,2/B],\dots,I_{B}=((B-1)/B,B]$.
Since the ECE is an estimator of TCE, it is important to understand the bias defined as
$\mathrm{Bias}(f_{w},S_{e},\mathrm{TCE})\coloneqq|\mathrm{TCE}(f_{w})-\mathrm{ECE}(f_{w},S_{e})|,$ (3)
and we refer to this as the total bias. As discussed in Section 1, one of our goals is to evaluate this bias theoretically in the multiclass classification setting.
2.3 Postprocessing by Recalibration
The trained predictor $f_{w}(x)$ can be poorly calibrated (Guo et al., 2017; Kumar et al., 2019). One common approach to address this problem involves recalibrating $f_{w}(x)$ by postprocessing using the parametric function $\eta_{V}:\Delta^{K}\to\Delta^{K}$, where $V\in\mathcal{V}\subset\mathbb{R}^{d^{\prime}}$ is the parameter learned by the recalibration dataset $S_{\mathrm{re}}\sim\mathcal{D}^{n_{\mathrm{re}}}$ at fixed $w$ and $S_{\mathrm{re}}$ is independent of $S_{\mathrm{tr}}$. In this procedure, the overall dataset is split into the training data $S_{\mathrm{tr}}$ for learning $W$, the recalibration data $S_{\mathrm{re}}$ for learning $V$, and the test data $S_{\mathrm{te}}$ for the evaluation of the ECE. We define the recalibrated model as $\eta_{v}\circ f_{w}$.
The output after recalibration, $\eta_{v}\circ f_{w}$, is expected to yield a sufficiently small $\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})$. From a generalization perspective, it is important to theoretically investigate conditions under which $\eta_{v}\circ f_{w}$ also achieves low $\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{te}})$. To this end, we define the following error term, resembling the standard generalization error typically defined via a loss function.
$\mathrm{gen}(\eta_{v}\circ f_{w},S_{\mathrm{re}},\mathrm{ECE})$ (4)
$\coloneqq|\mathbb{E}_{S_{\mathrm{te}}}\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{te}})-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})|.$ (5)
We refer to this quantity as the generalization error of the ECE under $S_{\mathrm{re}}$. In addition, the total bias for the recalibration, defined below, is required to be small:
$\mathrm{Bias}(\eta_{v}\circ f_{w},S_{\mathrm{re}},\mathrm{TCE})$ (6)
$\coloneqq|\mathrm{TCE}(\eta_{v}\circ f_{w})-\mathrm{ECE}(\eta_{v}\circ f_{w},S_{\mathrm{re}})|.$ (7)
Another goal of our study is to provide a theoretical understanding of these errors and biases by deriving their upper bounds using PAC-Bayes theory.
3 Analysis of ECE for Multiclass Classification
Here, we present a comprehensive analysis of the ECE before applying recalibration.
3.1 Total Bias Analysis for the ECE
In this section, we present our main analysis for the total bias. Under the smoothness assumptions commonly used in nonparametric estimation (Tsybakov, 2008), we show the following theorem:
Assumption 1.
Conditioned on $W=w$, the $L$-Lipschitz continuity holds for $P(Y=C|f_{w}(X)_{C})$.
@@ 3.1 Total Bias Analysis for the ECE
Given $W=w$, for any positive constant $\varepsilon\in(0,1)$ and $\lambda>0$, with probability $1-\varepsilon$ with respect to $S_{\mathrm{te}}$, we have
$\mathrm{Bias}(f_{w},S_{\mathrm{te}},\mathrm{TCE})\leq\frac{1+L}{B}+\frac{B\log 2+\log\frac{1}{\varepsilon}+\frac{2\lambda^{2}}{n_{\mathrm{te}}}}{\lambda}.$ (8)
The complete proof is shown in Appendix B.5. The first term on the right-hand side is referred to as the binning bias, whereas the second term is called the estimation bias (Futami and Fujisawa, 2024). According to this theorem, the number of bins shows a trade-off relationship as follows. As we increase $B$, the binning bias decreases because we estimate the label frequency more precisely. On the other hand, the statistical bias increases since the number of samples allocated to each bin decreases. From this observation, we derive the optimal number of bins that minimize the total bias. By setting $\lambda=\sqrt{Bn_{\mathrm{te}}}$, we can minimize the upper bound of Eq. (11) with respect to $B$. This yields $B=\mathcal{O}(n_{\mathrm{te}}^{1/3})$, and under this bin size, the order of the estimation bias becomes
$\mathrm{Bias}(f_{w},S_{\mathrm{te}},\mathrm{TCE})=\mathcal{O}(1/n_{\mathrm{te}}^{1/3}).$ (9)
This optimal order matches the result of the binary classification provided by Futami and Fujisawa (2024). This is natural, given that the binning ECE divides the top-$1$ predicted probabilities equally into $B$ bins in both cases.
The order of the total bias in Eq. (9) is tight from the nonparametric regression viewpoint. As discussed by Futami and Fujisawa (2024), the TCE measures the error between two functions, namely, $P(Y=C|f_{w}(x)_{C})$ and $f_{w}(x)$, and we estimate $P(Y=C|f_{w}(x)_{C})$ by binning, which is a nonparametric method. Thus, this is a problem of nonparametric regression on $[0,1]$ under Lipschitz continuity. According to Tsybakov (2008), the error in nonparametric regression cannot be smaller than $\mathcal{O}(1/n_{\mathrm{te}}^{1/3})$. Achieving an order smaller than this requires additional assumptions about the data distribution. Thus, the order of our bound is convincing under the current assumptions. We note that assuming Hölder continuity instead of Assumption 1 does not improve the order of the estimation bias owing to the bias caused by binning (see Appendix B.7 for details).
3.2 Total Bias Under the Training Dataset
Some recent algorithms (Kumar et al., 2018; Popordanoska et al., 2022; Wang et al., 2021a) focus on minimizing $\mathrm{ECE}(f_{w},S_{\mathrm{tr}})$ without guaranteeing that the gap between $\mathrm{TCE}(f_{w})$ and $\mathrm{ECE}(f_{w},S_{\mathrm{tr}})$ will remain small. Therefore, in this section, we theoretically investigate the conditions that an algorithm must satisfy to minimize this gap and to achieve a well-calibrated model. To this end, we conduct a theoretical analysis focusing on the following bias:
$\mathrm{Bias}(f_{w},S_{\mathrm{tr}},\mathrm{TCE})\coloneqq|\mathrm{TCE}(f_{w})-\mathrm{ECE}(f_{w},S_{\mathrm{tr}})|.$ (10)